# Sparse Repellency for Shielded Generation in Text-to-Image Diffusion Models

### v2

@@ Abstract
The increased adoption of diffusion models in text-to-image generation has triggered concerns on their reliability. Such models are now closely scrutinized under the lens of various metrics, notably calibration, fairness, or compute efficiency. We focus in this work on two issues that arise when deploying these models: a lack of diversity when prompting images, and a tendency to recreate images from the training set. To solve both problems, we propose a method that coaxes the sampled trajectories of pretrained diffusion models to land on images that fall outside of a reference set. We achieve this by adding repellency terms to the diffusion SDE throughout the generation trajectory, which are triggered whenever the path is expected to land too closely to an image in the shielded reference set. Our method is sparse in the sense that these repellency terms are zero and inactive most of the time, and even more so towards the end of the generation trajectory. Our method, named SPELL for sparse repellency, can be used either with a static reference set that contains protected images, or dynamically, by updating the set at each timestep with the expected images concurrently generated within a batch. We show that adding SPELL to popular diffusion models improves their diversity while impacting their FID only marginally, and performs comparatively better than other recent training-free diversity methods. We also demonstrate how SPELL can ensure a shielded generation away from a very large set of protected images by considering all 1.2M images from ImageNet as the protected set.
Stable Diffusion 3A close-up of an appleRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionSimple DiffusionThe Eiffel TowerRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionMDTv2ImageNet class 145Refer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to caption+ SPELL (Ours)Refer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to caption+ SPELL (Ours)Refer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to caption+ SPELL (Ours)Refer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to caption
Figure 1: SPELL interventions can change the diffusion trajectory of any pre-trained diffusion model, by self-avoiding other images generated within the same batch (here of size 6). SPELL achieves a higher intra-batch diversity using the same models, prompts and noise seeds.
@@ 1 Introduction
  • The SPELL mechanism adjusts diffusion trajectories when the generation of an image is expected to land close to a shielded region (defined as balls around the latent representation of images in the reference set). These regions can be static or updated dynamically, when used to obtain diversity, by using the expectation of other generation trajectories within the same batch.
  • SPELL interventions are sparse in that they only consider by design very few active shielded images (typically one) in the reference set at each time $t$. They are future looking in the sense that they are driven by geometric considerations on the expected final generation output at time 00, given position at time $t$, rather than on the position at time $t$ only as considered in (Corso et al., 2024; Zilberstein et al., 2024). As a result, they are mostly active at the start of the backward diffusion.
  • We show that applying SPELL to numerous state-of-the-art open-sourced diffusion models leads to images that better reflect the diversity of the true data (see Figure 1) with a marginal or even no increase in the Fréchet inception distance (FID) between geenrated images and training dataset.
@@
  • We showcase that SPELL can be scaled to a static reference set of millions of images, thanks to fast nearest-neigbor search tools. We use the whole ImageNet-1k dataset as the protected reference set, and generate images that are novel, without ever requiring to filter and/or regenerate images.
@@ 2 Background
Diffusion Models, also known as score-based generative models (Song et al., 2021; Song & Ermon, 2019; Ho et al., 2020), enable sampling from data distribution $p_{\textrm{data}}$ on support $\mathcal{X}\subset\mathbb{R}^{d}$, such as an image dataset, by simulating the reverse stochastic differential equation (SDE) (Haussmann & Pardoux, 1986; Anderson, 1965), initialised from some easy to sample prior $p_{1}\in\mathcal{P}(\mathbb{R}^{d})$:
$\mathrm{d}\mathbf{X}_{t}$ $=[f(t,\mathbf{X}_{t})-g^{2}(t)\nabla\log p_{t}(\mathbf{X}_{t})]\mathrm{d}t+g(t)\mathrm{d}B_{t}$ $\mathbf{X}_{1}$ $\sim p_{1},$ (1)
@@
Conditioning and Guidance. Conditional generation of diffusion generative models requires access to the conditional score $\nabla\log p_{t}(x_{t}\mid y)$ for some condition $y$ such as text or label. It is typically approximated either with explicit conditioning during training of the score / denoising network using equation 3 or as post-hoc additional guidance term added to the score. Given diffusion models have lengthy training procedures, likely due to their high variance loss (Jeha et al., 2024), it is desirable to guide diffusion models with inexpensive post-training guidance (Dhariwal & Nichol, 2021; Zhang et al., 2023; Denker et al., 2024), using e.g. classifier guidance (Dhariwal & Nichol, 2021)
@@
whereby the gradient $\nabla\log p(y\mid x_{t})$ of classifier $p(y\mid x_{t})$ for label $y$ is added to the score, heuristically multiplied by a scalar $\gamma\geq 1$ for increased guidance strength. Another approach which circumvents training a time-indexed classifier is using the approximation $p(y\mid x_{t})\approx{p}(y|\mathbf{X}_{0}=D_{\theta^{\star}}(t,x_{t}))$, for a pretrained denoiser $D_{\theta^{\star}}$ alias Diffusion Posterior Sampling (Chung et al., 2023).
Classifier-Free Guidance and Lack of Diversity in text-to-image Diffusion Models. Classifier-free guidance (CFG) (Ho & Salimans, 2022) is the dominant conditioning mechanism in text-to-image diffusion models, sharing properties with both explicit training and guidance. Similar to classifier guidance, CFG may be used to increase guidance strength but without resorting to approximating density $p(y\mid x_{t})$. Instead, the difference $\nabla\log p(y\mid x_{t})=\nabla\log p(x_{t}\mid y)-\nabla\log p_{t}(x_{t})$ is used as a guidance term, where each term is approximated with the same conditional network: $\nabla\log p(x_{t}\mid y)\approx s_{\theta}(t,x_{t},y)$ and $\nabla\log p(x_{t})\approx s_{\theta}(t,x_{t},\emptyset)$, for null condition $\emptyset$, trained as in equation 3. Adding CFG to the unconditional score yields $\nabla\log p_{t}(x_{t}\mid y)\approx\gamma s_{\theta}(t,x_{t},y)-(\gamma-1)s_{\theta}(t,x_{t},\emptyset)$. Despite its widespread popularity and good performance, CFG weighting is heuristic. It is not clear what final distribution is being generated; and practitioners observe a lack of diversity in generated samples (Somepalli et al., 2023; Chang et al., 2023; Wang et al., 2024b). In order to combat the lack of diversity from CFG, Corso et al. (2024) add to the score a repulsive interacting potential $\Phi$ evaluated on the whole batch, $(x^{(i)}_{t})_{i}$, coined particle guidance: $\nabla\log\Phi_{t}((x^{(i)}_{t})_{i})$. This repulsive potential is parameterised through a kernel as $\log\Phi_{t}((x^{(i)}_{t})_{i})=-\sum_{i,j}k_{t}(x_{t}^{(i)},x_{t}^{(j)})$. They specifically study the Gaussian kernel, Green’s function $k(x,y)\propto 1/\|x-y\|^{N}$ or kernels combined with a feature extractor $e_{\phi}$ such as DINO (Caron et al., 2021) $k_{t}(x,y)=k^{e}_{t}(e_{\phi}(x^{(i)}),e_{\phi}(y^{(j)}))$.
Figure 2: (a) At time $t$, we detect that the trajectory is headed (in expectation) into the shield of radius $r$ centered around $z_{2}$. Namely, we can compute $\mathbb{E}[\mathbf{X}_{0}\mid\mathbf{X}_{t}=x_{t}]$ using the pretrained denoiser evaluated on $x_{t}$, $D_{\theta^{\star}}(t,x_{t})$ to predict where the trajectory will land. Our sparse repellency (SPELL) term depicted as a black arrow adds a correction when generating $x_{t-\Delta t}^{\,}$ to ensure that the trajectory is pushed out of the shield. This is again in the case in the next step, when starting from $x_{t-\Delta t}$. (b) In batched generation, the shields / repellency balls are dynamically recreated at every iteration around each trajectory’s expected output. This prevents two elements in the batch, $x^{(1)}_{t}$ and $x^{(2)}_{t}$, from generating the same output thanks to corrections applied throughout their trajectory. (c) Both approaches can be combined to yield a diverse set of images, through intra-batch repellency, that won’t fall into protected images and/or previously generated images.
@@ 3 Sparse Repellency
We propose a mechanism to sample from the data density $p_{0}$ having the important requirement that generated samples $\mathbf{X}_{0}\sim p_{0}$ are far enough from each element of the reference set of repellency images $z_{i}\in\mathcal{X},k=1,\dots,K$. That set may be populated by real-world protected images, samples generated in earlier batches to increase diversity, images expected to be generated by other trajectories in the current batch, or a mix of all these types. More formally, we wish to sample a conditioned trajectory $\mathbf{X}_{t}\mid(\mathbf{X}_{0}\!\notin\!S)$, where $S$ is the collection of shields, i.e. balls of radius $r>0$ around repellency images, $S:=(\cup_{k}B_{k})$ with ${B_{k}=\{x\in\mathcal{X}\;:\;\|x-z_{k}\|_{2}\leq r\}}$. A blunt mechanism to guarantee generation outside of $S$ is to generate and discard: resample multiple times both initial noise and Brownian samples, follow the diffusion trajectory and repeat until a generated image falls outside of $S$. In the context of computationally expensive diffusion models, this would be wasteful and inefficient. We propose instead SPELL, a fairly simple approach that can be described from geometric principles and linked to DPS (Chung et al., 2023) and particle guidance Corso et al. (2024).
A Geometric Interpretation of SPELL. To ensure that generation falls outside of the shielded set $S$, our aim is to modify the diffusion trajectory at each time step, as presented in Figure 2, without having to discard any generation. To do so, we will rely at each time $t$ on the expected final output of the diffusion $\mathbb{E}[\mathbf{X}_{0}\mid\mathbf{X}_{t}=x_{t}]$ given current state $x_{t}$, as approximated by the diffusion network $D_{\theta^{\star}}(t,x_{t})$. We correct the trajectory whenever that expected vector falls within a shield. Using the notation $\hat{x}_{0}=D_{\theta^{\star}}(t,x_{t})$, we test whether for any index $k$ one has $\|\hat{x}_{0}-z_{k}\|_{2}<r$. If that is the case, the minimal modification $\delta$ that can ensure $\|\hat{x}_{0}+\delta-z_{k}\|_{2}\geq r$ is
@@
$\Delta=\sum_{k=1}^{K}\mathbf{1}_{\hat{x}_{0}\in B_{k}}\cdot\delta_{k}(\hat{x}_{0})=\sum_{k=1}^{K}\operatorname{ReLU}\left(\frac{r}{\|\hat{x}_{0}-z_{k}\|_{2}}-1\right)\cdot(\hat{x}_{0}-z_{k})\in\mathbb{R}^{d}$ (5)
@@
Figure 2(a) visualizes the repellence mechanism away from protected images while Figure 2(b) shows how it repels from trajectories within the same batch to enhance diversity. The batch generation produces $B$ samples $x^{(b)}_{0}$ in parallel and its repellency mechanism uses a time-evolving set of repellency points $z_{k,t}=\mathbb{E}[\mathbf{X}_{0}\mid\mathbf{X}_{t}=x^{(k)}_{t}]$ corresponding to the currently predicted end-state of each sample in the batch. As we made no further assumptions on $z_{k}$, these mechanisms can be mixed as in Figure 2(c) to enable diverse generation across arbitrary numbers of batches. This makes it possible to generate large numbers of diverse images even when the VRAM for each batch is limited. Note that SPELL is a post-hoc method that does not require retraining and can be applied to any diffusion score, in RGB space or latent space, unguided or classifier-free guided. Appendix D provides pseudo code and further implementation details.
SPELL as a DPS guidance term. We propose to derive SPELL as a guidance mechanism that can be tightly related to Doob’s $h$-transform. By Bayes’ rule
${\nabla_{x_{t}}\log p_{t}(x_{t}\mid x_{0}\notin S)=\nabla_{x_{t}}\log p_{t}(x_{t})+\nabla_{x_{t}}\log p_{0|t}(x_{0}\notin S\mid x_{t})}.$
@@
$\mathrm{d}\mathbf{X}_{t}=\left[f(t,\mathbf{X}_{t})-g(t)^{2}(\nabla\log p_{t}(\mathbf{X}_{t})+\nabla\log p_{0|t}(\mathbf{X}_{0}\notin S\mid x_{t}))\right]\mathrm{d}t+g(t)\mathrm{d}B_{t}.$ (6)
@@
$\nabla\log p_{0|t}(\mathbf{X}_{0}\notin S\mid x_{t})\approx\sum_{k=1}^{K}\omega(||\hat{x}_{0}-z_{k}||_{2},r)\cdot(\hat{x}_{0}-z_{k}),$ (7)
@@
(Intra-batch) SPELL as Particle Guidance. When used to promote diversity within the generation of a single batch, SPELL can be related to the self-interacting principle guidance (PG) proposed by Corso et al. (2024). That approach defines an interacting energy potential $\phi_{t}$ at time $t$, using the locations in space of all $B$ particles within a batch at time $t$. The gradient of that potential w.r.t. each particle, $\Delta^{(i)}=\nabla_{x^{(i)}_{t}}\log\Phi_{t}(x^{(1)}_{t},\ldots,x^{(B)}_{t})$ is then used to correct each individual trajectory to guarantee diversity. In contrast to this approach, SPELL draws insight on the expected future locations of points, at the end of the trajectory, i.e on the expected denoised images $\hat{x}_{0}^{(i)}$ and $\hat{x}_{0}^{(k)}$, where $\hat{x}_{0}=D_{\theta^{\star}}(t,x_{t})$. Indeed, the correction for each particle is explicitly given as:
$\Delta^{(i)}:=\sum_{k=1}^{K}\operatorname{ReLU}\left(\frac{r}{\|\hat{x}_{0}^{(i)}-\hat{x}_{0}^{(k)}\|_{2}}-1\right)\cdot(\hat{x}_{0}^{(i)}-\hat{x}_{0}^{(k)}).$ (8)
It is not clear yet whether the $B$ correction terms $\Delta^{(i)}$ considered by SPELL can be seen as the gradients of an interacting potential. While we can prove that $h(x)=\operatorname{ReLU}(\frac{r}{\|x\|}-1)x$ is indeed a conservative field (i.e. the gradient of a potential) as proved in Appendix C we do not know how to guarantee the same result for SPELL updates because of the composition of $h$ with the denoiser $D_{\theta^{\star}}$.
@@
4 Experiments
We now show that SPELL increases the diversity of modern text-to-image and class-to-image diffusion models (Section 4.2), with a better trade-off than other recent diversity methods (Section 4.3). We quantify the sparsity of SPELL interventions in Section 4.4 and utilize it to scale up to using all 1.2 million ImageNet-1k train images as a shielded set in Section 4.6.
4.1 Experimental Setup
We add SPELL to both class-to-image and text-to-image diffusion models. In the class-to-image setup, we use Masked Diffusion Transformers (MDTv2) (Gao et al., 2023), EDMv2 (Karras et al., 2024), and Stable Diffusion 3 Medium (SD3) (Esser et al., 2024). Since SD3 is a text-to-image model, we follow their template “a photo of a class_name”. We use the pretrained model checkpoints to generate 50,000 256x256 images of ImageNet-1k classes(Deng et al., 2009) without and with SPELL and compare them to the original ImageNet-1k images. We use the validation dataset as a comparison, since we will conduct experiments that repel from the training dataset in Section 4.6, which would render comparisons to the training dataset meaningless.
In our text-to-image setup, we use SD3, Latent Diffusion (Rombach et al., 2022), and RGB-space Simple Diffusion (Song et al., 2021) in resolution 256x256. For the latter two, we use the checkpoints of Gu et al. (2023). Details on hyperparameters are provided in Appendix D. We evaluate these models on CC12M (Changpinyo et al., 2021). As we target the ground truth diversity of images related to each prompt, we select a subset of captions with multiple corresponding images. This gives a one-to-many setup with 5000 captions and 4 to 128 images each (in total 41,596 images). We explain the construction of this dataset in Appendix E.
To evaluate diversity, we track the recall (Kynkäänniemi et al., 2019), coverage (Naeem et al., 2020), and Vendi score (Friedman & Dieng, 2023). To evaluate image quality, we use precision (Kynkäänniemi et al., 2019) and density (Naeem et al., 2020). We track these metrics per class/prompt and average across classes/prompts. To measure whether the generated images match the true image distributions, we use the marginal Fréchet Inception Distance (FID) (Heusel et al., 2017) and the more recent marginal Fréchet Distance with DINOv2 features ($\text{FD}_{\text{DINOv2}}$, Stein et al. (2024); Oquab et al. (2024)).
4.2 Benchmark

### v3

# Shielded Diffusion: Generating Novel and Diverse Images using Sparse Repellency

@@ Abstract
The adoption of text-to-image diffusion models raises concerns over reliability, drawing scrutiny under the lens of various metrics like calibration, fairness, or compute efficiency. We focus in this work on two issues that arise when deploying these models: a lack of diversity when prompting images, and a tendency to recreate images from the training set. To solve both problems, we propose a method that coaxes the sampled trajectories of pretrained diffusion models to land on images that fall outside of a reference set. We achieve this by adding repellency terms to the diffusion SDE throughout the generation trajectory, which are triggered whenever the path is expected to land too closely to an image in the shielded reference set. Our method is sparse in the sense that these repellency terms are zero and inactive most of the time, and even more so towards the end of the generation trajectory. Our method, named SPELL for sparse repellency, can be used either with a static reference set that contains protected images, or dynamically, by updating the set at each timestep with the expected images concurrently generated within a batch, and with the images of previously generated batches. We show that adding SPELL to popular diffusion models improves their diversity while impacting their FID only marginally, and performs comparatively better than other recent training-free diversity methods. We also demonstrate how SPELL can ensure a shielded generation away from a very large set of protected images by considering all 1.2M images from ImageNet as the protected set.
Machine Learning, ICML
Stable Diffusion 3Prompt: “A close-up of an apple”Refer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionSimple DiffusionPrompt: “The Eiffel Tower”Refer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionMDTv2ImageNet class 145 (king penguin)Refer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to caption+ SPELL (Ours)Refer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to caption+ SPELL (Ours)Refer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to caption+ SPELL (Ours)Refer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to caption
Figure 1: SPELL interventions can change the diffusion trajectory of any pre-trained diffusion model by self-avoiding other images generated, in the same or previous batches (and also any other non-generated image). This makes SPELL achieve a higher diversity above, with prompts, and noise seeds as the base models. We provide more qualitative examples in Appendix N.
@@ 1 Introduction
  • SPELL interventions are sparse by design; they only consider very few active shielded images (typically one) at each time $t$, and happen mostly early in the generation.
  • SPELL can increase the diversity of outputs by dynamically updating the shielded reference set to be images that were generated in previous batches and those that are expected to be generated in the current batch.
  • We apply SPELL to numerous state-of-the-art diffusion models and find that the generated images better reflect the diversity of the true data (see Figure 1) with a marginal or even no increase in the Fréchet inception distance (FID).
@@
  • We scale SPELL to a reference set of millions of images. This enables a second use-case: We shield the whole ImageNet-1k train dataset and generate images that are novel, without requiring to filter or regenerate images.
@@ 2 Background
Diffusion Models, also known as score-based generative models (Song et al., 2021; Song & Ermon, 2019; Ho et al., 2020), enable sampling from data distribution $p_{\textrm{data}}$ on support $\mathcal{X}\subset\mathbb{R}^{d}$, such as an image dataset, by simulating the reverse stochastic differential equation (SDE) (Haussmann & Pardoux, 1986; Anderson, 1965), initialised from some easy to sample prior $p_{1}\in\mathcal{P}(\mathbb{R}^{d})$, $\mathbf{X}_{1}\sim p_{1}$:
$\mathrm{d}\mathbf{X}_{t}=[f(t,\mathbf{X}_{t})-g^{2}(t)\nabla\log p_{t}(\mathbf{X}_{t})]\mathrm{d}t+g(t)\mathrm{d}B_{t},$ (1)
@@
Conditioning and Guidance. Conditional diffusion models requires access to the conditional score $\nabla\log p_{t}(x_{t}\mid y)$ for some condition $y$ such as text or label. It is typically approximated either with explicit conditioning during training of the score / denoising network using equation 3 or as post-hoc additional guidance term added to the score. Given diffusion models have lengthy training procedures, likely due to their high variance loss (Jeha et al., 2024), it is desirable to guide diffusion models with inexpensive post-training guidance (Dhariwal & Nichol, 2021; Zhang et al., 2023; Denker et al., 2024), using e.g. classifier guidance (Dhariwal & Nichol, 2021)
@@
whereby the gradient $\nabla\log p(y\mid x_{t})$ of classifier $p(y\mid x_{t})$ for label $y$ is added to the score, heuristically multiplied by a scalar $\gamma\geq 1$ for increased guidance strength. Another approach which circumvents training a time-indexed classifier is using the approximation $p(y\mid x_{t})\approx{p}(y|\mathbf{X}_{0}=D_{\theta^{\star}}(t,x_{t}))$, for a pretrained denoiser $D_{\theta^{\star}}$ alias Diffusion Posterior Sampling (DPS) (Chung et al., 2023).
Classifier-Free Guidance and Lack of Diversity in text-to-image Diffusion Models. Classifier-free guidance (CFG) (Ho & Salimans, 2022) is the dominant conditioning mechanism in text-to-image diffusion models, sharing properties with both explicit training and guidance. Similar to classifier guidance, CFG may be used to increase guidance strength but without resorting to approximating density $p(y\mid x_{t})$. Instead, the difference $\nabla\log p(y\mid x_{t})=\nabla\log p(x_{t}\mid y)-\nabla\log p_{t}(x_{t})$ is used as a guidance term, where each term is approximated with the same conditional network: $\nabla\log p(x_{t}\mid y)\approx s_{\theta}(t,x_{t},y)$ and $\nabla\log p(x_{t})\approx s_{\theta}(t,x_{t},\emptyset)$, for null condition $\emptyset$, trained as in equation 3. Adding CFG to the unconditional score yields $\nabla\log p_{t}(x_{t}\mid y)\approx\gamma s_{\theta}(t,x_{t},y)-(\gamma-1)s_{\theta}(t,x_{t},\emptyset)$. Despite its widespread popularity and good performance, CFG weighting is heuristic. It is not clear what final distribution is being generated; and practitioners observe a lack of diversity in generated samples (Somepalli et al., 2023; Chang et al., 2023; Wang et al., 2024).
Figure 2: (a) At time $t$, by computing $\mathbb{E}[\mathbf{X}_{0}\mid\mathbf{X}_{t}=x_{t}]$, we detect that the trajectory is headed (in expectation) into the shield of radius $r$ centered around $z_{2}$. Our sparse repellency (SPELL) term depicted as a black arrow adds a correction when generating $x_{t-\Delta t}^{\,}$ to ensure that the trajectory is pushed out of the shield. This is again in the case in the next step, when starting from $x_{t-\Delta t}$. (b) In batched generation, the shields are dynamically recreated at every iteration around each trajectory’s expected output. This prevents two elements in the batch, $x^{(1)}_{t}$ and $x^{(2)}_{t}$, from generating the same output. (c) Both approaches can be combined to yield a diverse set of images that won’t fall into protected images and previously or concurrently generated images.
@@ 3 Sparse Repellency
In this section, we introduce SPELL. We first give a geometrical intuition of how its repellency terms steer the diffusion trajectory out of shielded areas. Then, we make a deeper dive into theoretical connections to DPS (Chung et al., 2023) and particle guidance (Corso et al., 2024).
Setup. Our goal is to sample from the data distribution $p_{0}$ whilst satisfying the important requirement that generated samples $\mathbf{X}_{0}\sim p_{0}$ are far enough away from each element of the reference set of repellency images $z_{i}\in\mathcal{X},k=1,\dots,K$. That set may be populated by real-world protected images, samples generated in earlier batches, images expected to be generated by other trajectories in the current batch, or a mix of all these types. More formally, we wish to sample a conditioned trajectory $\mathbf{X}_{t}\mid(\mathbf{X}_{0}\!\notin\!S)$, where $S$ is the collection of shields, i.e. balls of radius $r>0$ around repellency images, $S:=(\cup_{k}B_{k})$ with ${B_{k}=\{x\in\mathcal{X}\;:\;\|x-z_{k}\|_{2}\leq r\}}$. A brute-force mechanism to guarantee generation outside of $S$ is to generate and discard: resample multiple times both initial noise and Brownian samples, follow the diffusion trajectory and repeat until a generated image falls outside of $S$. In the context of computationally expensive diffusion models, this would be wasteful and inefficient. Instead, we seek a mechanism which satisfies the protection in each generation, for any conditional or unconditional diffusion model.
A Geometric Interpretation of SPELL. To ensure that generation falls outside of the shielded set $S$, we modify the diffusion trajectory at each time step, as presented in Figure 2, without having to discard any samples. To do so, we correct the trajectory whenever the expected output, $\mathbb{E}[\mathbf{X}_{0}\mid\mathbf{X}_{t}=x_{t}]$ falls within a shield. Here the expected given current state $x_{t}$, is approximated by the denoising network $D_{\theta^{\star}}(t,x_{t})$. Using the notation $\hat{x}_{0}=D_{\theta^{\star}}(t,x_{t})$, we test whether for any index $k$ one has $\|\hat{x}_{0}-z_{k}\|_{2}<r$. If that is the case, the minimal modification $\delta$ that can ensure $\|\hat{x}_{0}+\delta-z_{k}\|_{2}\geq r$ is
@@
$\Delta$ $=\sum_{k=1}^{K}\mathbf{1}_{B_{k}}(\hat{x}_{0})\cdot\delta_{k}(\hat{x}_{0})$ (5)
$=\sum_{k=1}^{K}\operatorname{\tiny\sigma_{\tiny relu}}\left(\frac{r}{\|\hat{x}_{0}-z_{k}\|_{2}}-1\right)\cdot(\hat{x}_{0}-z_{k})\in\mathbb{R}^{d}$
@@
Figure 2(a) visualizes the repellence mechanism away from static protected images while Figure 2(b) shows how it dynamically repels from trajectories within the same batch. The batch generation produces $B$ samples $x^{(b)}_{0}$ in parallel and its repellency mechanism uses a time-evolving set of repellency points $z_{k,t}=\mathbb{E}[\mathbf{X}_{0}\mid\mathbf{X}_{t}=x^{(k)}_{t}]$ corresponding to the currently predicted end-state of each sample in the batch. Overall, this leads to a blue-noise like coverage of the distribution, as we visualize in 2D for the two-moons dataset in Figure 3. As we made no further assumptions on $z_{k}$, these mechanisms can be mixed as in Figure 2(c) to enable diverse generation across arbitrary numbers of batches. This makes it possible to generate large numbers of diverse images even when the VRAM for each batch is limited. Note that SPELL is a post-hoc method that does not require retraining and can be applied to any diffusion score, in RGB space or latent space, unguided or classifier-free guided. Appendix D provides pseudo code and further implementation details.
SPELL as a DPS guidance term. We propose to derive SPELL as a guidance mechanism, by Bayes’ rule
$\nabla_{x_{t}}\log p_{t}(x_{t}\mid x_{0}\notin S)=$ $\nabla_{x_{t}}\log p_{t}(x_{t})$
$+\nabla_{x_{t}}\log p_{0|t}(x_{0}\notin S\mid x_{t})$
@@
$\tilde{s}_{t}(\mathbf{X}_{t},S)=\nabla\log p_{t}(\mathbf{X}_{t})+\nabla\log p_{0|t}(\mathbf{X}_{0}\notin S\mid x_{t})$
$\mathrm{d}\mathbf{X}_{t}=\left[f(t,\mathbf{X}_{t})-g(t)^{2}\tilde{s}_{t}(\mathbf{X}_{t},S)\right]\mathrm{d}t+g(t)\mathrm{d}B_{t}.$ (6)
@@
$\nabla\log p_{0|t}(\mathbf{X}_{0}\notin S\mid x_{t})\approx\sum_{k=1}^{K}\omega(||\hat{x}_{0}-z_{k}||_{2},r)\cdot(\hat{x}_{0}-z_{k}),$
@@
(a) DDPM without SPELL
(b) DDPM + Intra-batch SPELL ($r=0.075$)
Figure 3: 200 samples generated with an unconditional DDPM diffusion model on the two moons toy dataset, with and without intra-batch SPELL with minimum shield radius of $r=0.075$. SPELL’s shields lead to a blue-noise-like coverage of the distribution.
(Intra-batch) SPELL and Particle Guidance. When using SPELL to promote diversity within the generation of a single batch (but without the more general protection against arbitary or previously generated images), SPELL can be related to the self-interacting particle guidance (PG) approach proposed by (Corso et al., 2024). That approach defines an interacting energy potential $\phi_{t}$ at time $t$, using the locations in space of all $B$ particles within a batch at time $t$. The gradient of that potential w.r.t. each particle, $\Delta^{(i)}=\nabla_{x^{(i)}_{t}}\log\Phi_{t}(x^{(1)}_{t},\ldots,x^{(B)}_{t})$ is then used to correct each individual trajectory to guarantee diversity. In contrast to this approach, SPELL draws insight on the expected future locations of points, at the end of the trajectory, i.e on the expected denoised images $\hat{x}_{0}^{(i)}$ and $\hat{x}_{0}^{(k)}$, where $\hat{x}_{0}=D_{\theta^{\star}}(t,x_{t})$. Indeed, the correction for each particle is explicitly given as:
$\Delta^{(i)}:=\sum_{k=1}^{K}\operatorname{\tiny\sigma_{\tiny relu}}\left(\frac{r}{\|\hat{x}_{0}^{(i)}-\hat{x}_{0}^{(k)}\|_{2}}-1\right)\cdot(\hat{x}_{0}^{(i)}-\hat{x}_{0}^{(k)}).$ (7)
The $B$ correction terms $\Delta^{(i)}$ considered by SPELL cannot be seen to our knowledge as the gradients of an interacting potential. While we prove in Appendix C that $h(x)=\operatorname{\tiny\sigma_{\tiny relu}}(\frac{r}{\|x\|}-1)x$ is a conservative field (i.e. the gradient of a potential), we find no guarantee for the more complex SPELL updates above which involve compositions of $h$ with the denoiser $D_{\theta^{\star}}$. Even if SPELL was a conservative field, the biggest difference between PG and SPELL is that PG defines dense interventions between all particles using soft-vanishing kernels that are never zero and thus always perturb diffusion trajectories. SPELL, conversely, intervenes sparsely and rarely, both in time and w.r.t. points in the reference set. As a result, the original diffusion process is less perturbed, notably towards the end of a trajectory, and SPELL scales to large reference sets of millions of shields.
@@
4 Related Works
Most closely related to our method is that of Corso et al. (2024), who promote diversity through an intra-batch repulsion term, used as guidance for pre-trained diffusion models. Similarly to our work the repulsion term can be applied at $\mathbb{E}[\mathbf{X}_{0}|x_{t}]$ or on features. The primary difference is the sparsity of the repulsion term in SPELL, and using this on a fixed reference set in addition to intra-batch.